1 number is 4 less than 3 times another. If their sum is increased by 5, the result is 25. Find the numbers.
The numbers are 6 & 14 - how did they get 6, I can figure out 14 but missing an essential part.
Thank you.

Found 2 solutions by craigt, meow123:
Answer by craigt(10)

Two numbers, x & y
y=3x-4
x + y + 5 = 25
x + 3x - 4 + 5 = 25
4x +1 = 25
4x = 24
x = 6, So one of the numbers is 6, now solve for y
y = 3x-4 = 3(6)-4 = 14

Answer by meow123(2)

two donkeys and black jaguar